OpenStudy (anonymous):

Determine the number of atoms in each sample. a.) 98.3 g mercury, Hk c.) 10.7 g lithium, Li Determine the number of moles in each sample. a.) 6.84g sucrose C12H22)11 c.) 68.0g ammonia, NH3 Determine the mass of the following molar quantity. a.) 3.52 mol Si c.) 0.550 mol F2 Chemistry, year1, honors. Can *someone* PLEASE help me?

OpenStudy (anonymous):

1 mole = 6.02 x 10^23 atoms, avogrado constnt to find how much moles the substances, use the formula: (relative mass)/(molar mass - u can find this in periodic table)= number of mole periodic table http://en.wikipedia.org/wiki/Periodic_table reminder ''1 mole = 6.02 x 10^23 atoms'' use the number of mole u get times the 6.02....to get the number of atoms in the substance. For example, mecury, Hg , not Hk 98.3g/80g=1.22875 moles number of atom = 1.22875 x 6.02 x 10^23=7.397075 x 10^23 try the rest or u will not learn it, good luck, hope u understand ^^

OpenStudy (anonymous):

if it is a compound, like NH3, u neeed to sum up all the component in it N=7g h=1g , NH3 has 1N and 3 H = 7g+1g +1g +1g = 10g(molar masss of NH3) from mole to mass: mass of n mole of substance = n mole x molar mass like Si, silicon: 3.52 mole silicon 's mass = 3.52 x 14g =??g

OpenStudy (anonymous):

Thankyou! :)

OpenStudy (preetha):

Samana, if you show your work here, we can help you if you go wrong. That way you'll do it right.

OpenStudy (preetha):

Here are the relationships: g to atom: Number of atoms = Mass of element in g x (1 mole/Mass of 1 mole (g)) x( 6.023 x 10 23 atoms/1mole)

OpenStudy (preetha):

So g and moles cancel out and leave atoms

OpenStudy (preetha):

You should always set up your work like this and make sure the units cancel out to leave you what you want. Also always give your answer with the correct units, g or atoms or moles or what ever.